1.solve the following differential equation by using separation of variables method:

1.1.x(dy/dx)=4y.

2.show whether or not the following differential equations are separable:

2.1.dy/dx=t(in(s^2t))+8t².

2.2.dy/dx=ye^x+y/x²+2.

2.3.dy/dx=x+1/y-1.


1
Expert's answer
2021-07-09T05:13:04-0400

Question 1:

Solve the following differential equation by using seperation of variables method:

1.1 x"\frac{dy}{dx}=4y"

Solution:

The equation can be written as

"\frac{dy}{4y}=\frac{dx}{x}"

"\int\frac{dy}{4y}=\int\frac{dx}{x}"

"\frac{ln(|y|)}{4}={ln(|x|)}+C"

Answer

"\frac{ln(|y|)}{4}=ln(|x|)+C"

Question 2

Show whether or not the following differential equations are separable.

2.1.dy/dx=t(in(s^2t))+8t2

2.2.dy/dx=ye^x+y/x2+2

2.3.dy/dx=x+1/y-1

Solution;

2.1

"\frac {dy}{dx}=t(in(s^2t))+8t^2"

The equation is inseparable.

It's a parametric equations. The variables of integration do not match those of the equation.

2.2

"\frac{dy}{dx}=ye^x+\frac{y}{x^2}+2"

"\frac{dy}{dx}=y(e^x+\frac{1}{x^2})+2"

"\frac{dy}{y}=((e^x+\frac{1}{x^2})+\frac{2}{y}))dx"

The equation is inseparable.

2.3

"\frac{dy}{dx}=\frac{x+1}{y-1}"

(y-1)dy=(x+1)dx

The equation is separable.